OpenStudy (zasharra):

f(x) = 3x + 9, g(x) = 5x^2 Find (f + g)(x).

OpenStudy (zasharra):

@.Sam. @sweetburger @zepdrix

OpenStudy (zasharra):

@kidrah69

OpenStudy (zasharra):

@zepdrix do you know what to do?

zepdrix (zepdrix):

\[\large\rm \color{royalblue}{f=3x+9}\qquad\qquad \color{orangered}{g=5x^2}\]Should be pretty straight forward,\[\large\rm \color{royalblue}{f}+\color{orangered}{g}=\color{royalblue}{3x+9}+\color{orangered}{5x^2}\]ya? :d

OpenStudy (zasharra):

that's what i thought but this is my first time doing preC so i wnted to make sure I have a few more you down?

zepdrix (zepdrix):

sure let's try another :D

OpenStudy (zasharra):

f(x) = 2x + 6, g(x) = 4x2 Find (f + g)(x).

OpenStudy (zasharra):

2x+6/4x^2 is it?

zepdrix (zepdrix):

Why division? :o (f+g)(x) basically just means add f and g.

OpenStudy (zasharra):

that's a fraction

zepdrix (zepdrix):

fraction = division

OpenStudy (zasharra):



OpenStudy (zasharra):

is that the correct answer ?

OpenStudy (zasharra):

Question: f(x) = 2x + 6, g(x) = 4x2 Find (f + g)(x).

zepdrix (zepdrix):

I don't understand why you would guess the fraction... Our first problem was almost exactly like this one :o We didn't end up with a fraction. We just added the pieces together.

OpenStudy (zasharra):

yes it did? what these are my answer choices to the first problem: -5x2 + 3x + 9 3x+9/5x^2 15x3 + 45x 5x2 + 3x + 9

OpenStudy (zasharra):

B was the answer right?

OpenStudy (zasharra):

it was D wasn't it

zepdrix (zepdrix):

f=3x+9 g=5x^2 So f+g is going to be 3x+9+5x^2 which no, is not equivalent to B.

zepdrix (zepdrix):

Yes D. Just rearrange 3x+9+5x^2 = 5x^2+3x+9

zepdrix (zepdrix):

I didn't put a fraction bar when I wrote the answer to the first one. Not sure where you got that from XD hehe

OpenStudy (zasharra):

f(x) = 2x + 6, g(x) = 4x2 Find (f + g)(x). (1 points) a. 8x3 + 24x b. 4x2 + 2x + 6 c. -4x2 + 2x + 6

OpenStudy (zasharra):

b?

zepdrix (zepdrix):

Mmm yes, good. These pieces added together.

OpenStudy (zasharra):

f(x) = 4x + 6, g(x) = 2x2 Find (fg)(x). a. 8x3 + 12x2 b. 2x2 + 4x + 6 c. 8x + 12 d. 8x2 + 12x

OpenStudy (zasharra):

b again?